7 January 2014 – Tuesday

LDAP:

  • Four issues – Three former students – did the dance three times.  One employee – that one might be a “former” graduate student and I sent a query off to HR to confirm.
  • Checking up on the changes to the development server for the ou and title attributes – no problems with last night’s update.
  • Installed the ldif_diff code change (not the updated list of replace only attributes) from the development system onto the production ldap server.

Backups:

  • 40 tapes to the vaults, 40 tapes from the vaults.
  • E-Mail backups were messed up.  Kent rebooted the world and re-ran the snap. I restarted the backups. It is assumed this is fallout from last night’s SAN upgrade.
  • Clone status (what’s left):
    • Tape: 1603.6 GB
    • COMIS: 7118.9 GB
    • Disk: 41714.1 GB
      • That disk one may not finish, but the other two should!

Tags:

Leave a Reply

You must be logged in to post a comment.